A ladder is leaning against a building. The distance from the bottom of the ladder to the building is 8 ft less then the length of the ladder. How high up the side of the building is the top of the ladder if that distance is 4 ft less then the length of the ladder?
Answer by RAY100(1637)

assume problem is right triangle with sides (l-8), (l-4), and hypotenuse of l
pythagorous theorem states c^2 = a^2 + b^2
l^2 = (l-8)^2 + (l-4)^2
l^2 = ( l^2 -16 * l +64) + (l^2 -8*l +16)
L^2 = 2*l^2 - 24*l + 80
0 = l^2 -24*l + 80
factoring 0 = ( l -20) ( L-4)
l-20 = 0. l=+20
l-4 = 0, l=4 but l-8 and l-4 are positive lengths therefore l cannot equal 4
sides of triangle are 20, l-8 = 12, and l-4 =16 ( height up bldg)